OpenStudy (anonymous):

find the distance between the points (5,-8) and (10,-3). Also find the midpoint of the segment that joins the points

OpenStudy (anonymous):

the midpoint is easiest, take the average in each coordinate the average of \(5\) and \(10\) is \(\frac{10+5}{2}=\frac{15}{2}=7.5\)

OpenStudy (anonymous):

the average of \(-3\) and \(-8\) is \(\frac{-3-8}{2}=-\frac{11}{2}=-5.5\)

OpenStudy (anonymous):

so the midpoint is \[(7.5,-5.5)\]